OpenStudy (anonymous):

simplify the expression 19Y/y+2 - 7y/y^2-4

OpenStudy (lalaly):

\[(\frac{19y}{y+2} \times \frac{y-2}{y-2}) -\frac{7y}{(y-2)(y+2)}\]\[\frac{19y(y-2)-7y}{(y-2)(y+2)}\]

OpenStudy (lalaly):

\[\frac{19y^2-38y-7y}{(y-2)(y+2)}\]

OpenStudy (lalaly):

\[\frac{19y^2-45y}{(y-2)(y+2)}\]